OpenStudy (anonymous):

Will medal and fan!! Can someone explain to me how to do Quadratic Trinomials?

OpenStudy (anonymous):

@peak

OpenStudy (anonymous):

i suck at quadratic trinominals. sry

OpenStudy (anonymous):

hola

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

u espanic

OpenStudy (anonymous):

meda

OpenStudy (anonymous):

u espanic si

OpenStudy (anonymous):

i know a little

OpenStudy (anonymous):

And its okay @peak

OpenStudy (anonymous):

amigos

OpenStudy (anonymous):

lol its a prank im not rally Spanish but i know a little

OpenStudy (anonymous):

@Christian_10_

OpenStudy (anonymous):

i could tell xD

OpenStudy (anonymous):

how

OpenStudy (anonymous):

You weren't very good at speaking it and i believe you spelled a couple words wrong lol

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

lol im such a retart

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

and don't say that

OpenStudy (anonymous):

What kind of questions? Quadratic --> highest power of x is x^2, Trinomials ---> three terms added.

OpenStudy (anonymous):

One of my homework questions was 3r^2-6r-45

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Who keeps on giving me medals

OpenStudy (anonymous):

im stupid

OpenStudy (anonymous):

\[3r^2-6r-45\] it's got three term, and 2 is the highest power of r, so what do you need to do with it?

OpenStudy (anonymous):

I think its set up like this (3r-9)(3r-5)

OpenStudy (anonymous):

When you multi back out... 3r*3r = 9r^2. so that's not right. try and factor the three out of everything first. i.e. 3(---------------)

OpenStudy (anonymous):

3(r-9)(r-5)?

OpenStudy (anonymous):

im sorry im not good at math /.\

OpenStudy (anonymous):

So to factor out the 3, you want \[3*1r^2+3*2r+3*15=3(r^2+2r+15)\]

OpenStudy (anonymous):

opps, negatives! \[3(r^2-2r-15)\]

OpenStudy (anonymous):

so that's the answer?

OpenStudy (anonymous):

now you want \[(r+a)(r+b)\] such that a*b=-15 and a+b = -2.

OpenStudy (anonymous):

half way

OpenStudy (anonymous):

to find a*b = -15, you can find all the ways to multiply to get -15... -15*1 -5*3 -3*5 -1*15

OpenStudy (anonymous):

and then check if any give a+b=-2.... and -5+2 = -2!!

OpenStudy (anonymous):

putting it all together \[3(r-5)(r+2)\]

OpenStudy (anonymous):

hope that makes sense, gtg.

OpenStudy (anonymous):

thank youu!!!!!!

OpenStudy (anonymous):

np